# El Gordo needs El Anzuelo: Probing the structure of cluster members with multi-band extended arcs in JWST data

### v1

@@ A.1 Surface brightness properties of the deflectors
A.2 Full joint posterior distribution of mass model parameters
@@ A.3 Robustness of the constraints
A.4 Fifth and central images in strong lenses
As shown in Sect. 5.4, we confirm the five-image lensing configuration of El Anzuelo. This is a rather rare situation: in only a few galaxy-scale strong lenses with a single or two main deflectors has been discovered a fifth extended, non-demagnified image. The vast majority of fifth images fall in the category of unresolved demagnified central images, which appear close to the deflector centroid and are often only detected at radio wavelengths thanks to weaker contamination from the lens light (e.g., Gavazzi et al. 2003; Tamura et al. 2015; Wong et al. 2015; Collett et al. 2017). The appearance of extended source features in the vicinity of L1 and L3 centroids may provide constraints on the inner part of the mass density profile of the deflectors (at the cost of larger blending between the light of the various components). A comparable configuration in the literature is the Cosmic Horseshoe strong lens, for which Schuldt et al. (2019) used the image of a second lensed source appearing close to the lens galaxy to constrain the inner dark matter profile of that galaxy. However, as it is the case for the Cosmic Horseshoe, the fifth image often corresponds to a radial image (as opposed to much more common tangential images). Coming back to El Anzuelo, the fifth of the source remains a tangential one, thus making it a very peculiar lensing configuration. Although not entirely comparable, strong lens systems in which a low-mass satellite galaxy coincides with a strongly lensed arc and locally creates additional strong lensing features can give rise to additional tangential images (see systems analysed in Suyu & Halkola 2010; Vegetti et al. 2010).
Appendix B Spectral energy distribution of a source clump

### v2

@@ C.1 Surface brightness properties of the deflectors
C.2 Full joint posterior distribution of mass model parameters
@@ C.3 Robustness of the constraints
C.4 Fifth and central images in strong lenses
As shown in Sect. 5.4, we confirm the five-image lensing configuration of El Anzuelo. This is a rather rare situation: in only a few galaxy-scale strong lenses with a single or two main deflectors has been discovered a fifth extended, non-demagnified image. The vast majority of fifth images fall in the category of unresolved demagnified central images, which appear close to the deflector centroid and are often only detected at radio wavelengths thanks to weaker contamination from the lens light (e.g., Gavazzi et al. 2003; Tamura et al. 2015; Wong et al. 2015; Collett et al. 2017). The appearance of extended source features in the vicinity of L1 and L3 centroids may provide constraints on the inner part of the mass density profile of the deflectors (at the cost of larger blending between the light of the various components). A comparable configuration in the literature is the Cosmic Horseshoe strong lens, for which Schuldt et al. (2019) used the image of a second lensed source appearing close to the lens galaxy to constrain the inner dark matter profile of that galaxy. However, as it is the case for the Cosmic Horseshoe, the fifth image often corresponds to a “radial” image (as opposed to much more common “tangential” images). Coming back to El Anzuelo, the fifth of the source remains a tangential one, thus making it a very peculiar lensing configuration. Although not entirely comparable, strong lens systems in which a low-mass satellite galaxy coincides with a strongly lensed arc and locally creates additional strong lensing features can give rise to additional tangential images (see systems analysed in Suyu & Halkola 2010; Vegetti et al. 2010).
Appendix D Spectral energy distribution of a source clump